OpenStudy (anonymous):

Solve for y, then find the value of y when given x=-2 6x=7-4y

OpenStudy (the_fizicx99):

To solve for y, 6x = 7 -4y -7 -7 -7 + 6x = -4y /-4 /-4 /-4 1.75 -1.5x = y 1.75 -1.5(-2) = y 1.75 + 3 = y 4.75 = y